OpenStudy (anonymous):

Which value is a solution to j – 6 is greater than or equal to 1? The Choices are. A. 1 B. 4 C. 6 D. 8

OpenStudy (ivettef365):

try plugging in the numbers j-6>= 1 1-6>=1 -5>=1 not true keep trying

OpenStudy (anonymous):

j - 6 >= 1 j >= 1 + 6 j >= 7 it would have to be D because j is GREATER or equal to 7